# Faithfulness-Aware Uncertainty Quantification for Fact-Checking the Output of Retrieval Augmented Generation

### v2

@@ 2.3 Calibrating franq
Since the UQ methods $\text{UQ}_{\text{faith}}$ and $\text{UQ}_{\text{unfaith}}$ of equation (5) are not inherently probabilistic and may have different distributions, to avoid inconsistencies and miscalibration among various UQ measures, we calibrate their outputs using isotonic regression on training data Vashurin et al. (2025).
Formally, given a training dataset $\mathcal{D}=\{(u_{i},\text{fact}_{i})\}_{i=1}^{N}$ comprising pairs of UQ scores $u_{i}$ and corresponding binary factuality labels $\text{fact}_{i}$ for $N$ claims, we calibrate UQ scores by fitting a non-decreasing function $f\colon\mathbb{R}\rightarrow[0,1]$ through isotonic regression, minimizing the squared error:
@@
where $\mathcal{F}$ denotes the set of all non-decreasing functions mapping real numbers to probabilities in the interval $[0,1]$. During inference, the calibration function $\widehat{f}$ is applied to each UQ score to produce probabilistically meaningful output.
Since $\text{UQ}_{\text{faith}}$ and $\text{UQ}_{\text{unfaith}}$ represent probabilities under faithful and unfaithful conditions, respectively, we propose condition-specific calibration. This involves partitioning the training dataset $\mathcal{D}$ into two subsets: faithful claims $\mathcal{D}_{\text{faith}}$ and unfaithful claims $\mathcal{D}_{\text{unfaith}}$. Then, we calibrate $\text{UQ}_{\text{faith}}$ using the faithful subset $\mathcal{D}_{\text{faith}}$ and $\text{UQ}_{\text{unfaith}}$ using the unfaithful subset $\mathcal{D}_{\text{unfaith}}$.
We consider franq with condition-specific calibration as our primary method. To evaluate the impact of calibration, we additionally assess two variants: one without any calibration, and another in which both UQ methods are calibrated using the full training dataset $\mathcal{D}$. The calibration strategies are summarized as follows:
@@
    Calibrated scores. Both UQ methods are calibrated on the entire training dataset $\mathcal{D}$, disregarding claim faithfulness.
@@
    Condition-calibrated scores. Each UQ method is calibrated using a subset of the training data corresponding to the respective condition: $\text{UQ}_{\text{faith}}$ is calibrated using $\mathcal{D}_{\text{faith}}$, and $\text{UQ}_{\text{unfaith}}$ is calibrated using $\mathcal{D}_{\text{unfaith}}$.
@@ 3 Datasets for RAG Uncertainty Quantification
Existing datasets for studying RAG hallucinations have serious limitations (see discussion in Section 5.2). We argue that an effective dataset should reflect generation behavior in the presence of retrieval, capturing both factual errors and contextual misuse. To this end, we introduce a new dataset specifically designed for long-form generations, enabling fine-grained analysis of atomic claims.
Overall, we consider two QA settings: long-form, where questions require comprehensive answers composed of multiple claims; and short-form, where questions are answered concisely with single claims. For both QA settings, we employ Llama-3.2-3B-Instruct Grattafiori et al. (2024) and Falcon3-3B-Base Team (2024) as the LLMs.
Long-form QA dataset. Generating high-quality data for RAG hallucination detection remains challenging. Most benchmarks focus on faithfulness (i.e., entailment with retrieved content) rather than factuality. However, we argue that factuality is more critical in RAG applications, with faithfulness serving as a complementary perspective. Since automatic annotation lacks precision and manual annotation is costly, we adopt a hybrid approach to ensure data quality.
Our long-form QA dataset consists of 76 questions: 44 from RAGTruth Niu et al. (2024) filtered to include questions where Llama 3B provides at least two false claims, and 32 technical questions generated via GPT-4 (e.g., “How does solar power generate electricity?”).
@@
We extract claims and their token spans from model outputs using methods introduced in Wang et al. (2024b); Vashurin et al. (2025). First, GPT-4o extracts decontextualized atomic claims from the entire text paragraph through a dedicated prompt. Then, for each claim, a second prompt instructs GPT-4o to list the relevant words from the original text, which we map to token spans. Applying this procedure, we obtain 1,782 claims for Llama 3B and 1,548 claims for Falcon 3B. From these claims, we select 500 claims for training, reserving the remainder for testing.
@@
We annotate factuality and faithfulness using GPT-4o-search and manually verify claims classified as incorrect. For faithfulness annotation, each claim is automatically assigned one of three categories: faithful, unfaithful-contra, or unfaithful-neutral. These categories are binarized for evaluation: faithful is mapped to 1, and both unfaithful categories to 0, as unfaithful-contra comprises less than 5% of the data. Similarly, factuality annotations include True, False, and unverifiable. Only verifiable claims (True or False) are retained, with labels binarized. Further details on prompts and dataset statistics are available in Appendix B.
@@
Short-form QA datasets. For short-form QA, we evaluate franq on four widely used QA datasets: TriviaQA Joshi et al. (2017), SimpleQA Wei et al. (2024), Natural Questions Kwiatkowski et al. (2019), and PopQA Mallen et al. (2023). The output is treated as a single claim. From each dataset, we sample 200 questions for training and 1,000 for testing.
We evaluate factuality by comparing RAG outputs with gold-standard answers using GPT-4, following the approach in Wei et al. (2024).
RAG models. For each question, we select top-$k$ passages with Facebook/contriever-msmarco retriever and Wikipedia embeddings based on the 2018 English Wikipedia, to ensure accurate retrievals. For long-form QA, we set $k=3$, and for short-form QA, $k=5$.
@@ 4 Experiments
For all the experiments, we fix both the retrieval process and the white-box LLM, verifying the factual accuracy of the LLM-generated claims.
We test franq with 3 calibration strategies (see Section 2.3). First, we compare it with the unsupervised baselines introduced earlier. While the uncalibrated version of franq is unsupervised, the calibrated versions require training labels. Therefore, we further compare our method against the supervised XGBoost baseline Kuleshov and Liang (2015) trained using the outputs of unsupervised UQ methods as features.
Furthermore, in ablation studies, we investigate the impact of different franq components $P(\text{faithful})$, $\text{UQ}_{\text{faith}}$, and $\text{UQ}_{\text{unfaith}}$, as well as varying training sizes.
@@ 4.1 Experimental Setup
Method Llama 3b Falcon 3b
ROC-AUC $\uparrow$ PR-AUC $\uparrow$ PRR $\uparrow$ ROC-AUC $\uparrow$ PR-AUC $\uparrow$ PRR $\uparrow$
Max Claim Probability .497 .058 -.029 .678 .126 .258
P(True) .573 .117 .207 .591 .077 .170
Perplexity .477 .056 -.081 .636 .090 .165
Mean Token Entropy .562 .109 .115 .646 .130 .219
CCP .587 .085 .169 .641 .162 .181
MaxNLI .642 .109 .151 .669 .101 .287
AlignScore .616 .075 .108 .652 .104 .233
Parametric Knowledge .523 .064 .018 .536 .067 .029
XGBoost (all UQ features) .611 .124 .206 .616 .088 .198
XGBoost (franq features) .576 .111 .149 .550 .080 .086
franq no calibration .646 .100 .181 .692 .135 .362
franq calibrated .653 .103 .256 .602 .074 .090
franq condition-calibrated .641 .140 .223 .693 .173 .354
@@
UQ baselines. We group all UQ methods into 4 categories: (1) general baselines, (2) RAG-specific baselines, (3) XGBoost-based methods, and (4) three variants of proposed franq method, each employing a different calibration strategy.
General baselines. To implement baselines, we use the LM-Polygraph library Fadeeva et al. (2023). For short-form QA, we use 13 sequence-level methods; for long-form QA, we use 5 claim-level methods. A complete list of these methods is provided in Table 1. We also evaluate three baseline methods previously introduced as components of franq: AlignScore, MaxNLI, and Parametric Knowledge (see Section 2.2).
XGBoost methods. We include XGBoost models trained on factuality labels using two feature sets: (1) three components used in franq (AlignScore, $\text{UQ}_{\text{faith}}$, $\text{UQ}_{\text{unfaith}}$), and (2) all available unsupervised UQ methods (for long-form QA: 5 general + 3 RAG-specific baselines; for short-form QA: 13 general + 3 RAG-specific baselines).
FRANQ. Finally, we evaluate three franq variants with different calibration strategies for $\text{UQ}_{\text{faith}}$ and $\text{UQ}_{\text{unfaith}}$ (see Section 2.3): “no calibration”, “calibrated”, and “condition-calibrated”.
Metrics. Each UQ method provides factuality estimates, which we compare with binary gold-standard factuality labels using ROC-AUC and PR-AUC (false claims are designated as the positive class to emphasize the method’s effectiveness in identifying them). To further check rejection performance, we evaluate Prediction Rejection Ratio (PRR; Mallen et al. (2023)) with a maximum rejection threshold of 0.5. It quantifies how well the model can reject uncertain predictions while retaining accurate ones, providing a direct measure of the model’s ability to identify and prioritize reliable outputs.
@@ 4.2 Long-Form QA Results
For long-form QA, we evaluate each UQ method using three metrics: ROC-AUC, PR-AUC, and PRR, across two models: Llama 3B and Falcon 3B. These results are presented in Table 2.
For Llama 3B, the calibrated version of franq achieves the highest ROC-AUC and PRR, indicating strong overall performance. The condition-calibrated version of franq yields the best PR-AUC and ranks second in PRR. For Falcon 3B, the condition-calibrated franq variant performs best overall, with the highest ROC-AUC and PR-AUC, while the non-calibrated franq achieves the top PRR and second-best ROC-AUC.
Note that, except for franq calibrated on Falcon 3B, the proposed franq method and its two calibration extensions achieve leading results on both models. In the case of franq calibrated on Falcon 3B, one hypothesis is that Parametric Knowledge calibration over the entire dataset degrades its performance. This may be because Parametric Knowledge on Falcon 3B data does not approximate faithfulness well enough, leading isotonic regression calibration to further degrade the final output.
Method Llama-3b Falcon-3b
MeanRank $\downarrow$ MeanValue $\uparrow$ MeanRank $\downarrow$ MeanValue $\uparrow$
ROC-AUC PR-AUC PRR ROC-AUC PR-AUC PRR ROC-AUC PR-AUC PRR ROC-AUC PR-AUC PRR
Max Sequence Probability 10.50 12.50 11.50 .758 .558 .454 14.75 12.50 14.00 .617 .628 .256
CCP 13.00 13.00 13.00 .745 .551 .443 11.75 10.50 10.50 .637 .641 .304
Max Token Entropy 7.25 6.75 8.00 .774 .594 .481 13.75 14.25 15.00 .617 .613 .242
P(True) 20.25 20.00 20.00 .477 .292 -.011 20.00 20.00 20.00 .515 .529 .042
Lexical Similarity 8.50 10.50 8.25 .769 .564 .479 13.00 14.00 13.50 .635 .618 .277
Degree Matrix 6.25 3.75 5.75 .789 .629 .520 3.00 3.00 3.25 .724 .702 .464
Eccentricity 6.00 5.00 6.50 .787 .615 .515 6.25 8.75 7.25 .702 .662 .406
Sum of Eigenvalues 5.25 4.25 5.50 .791 .628 .518 3.00 3.50 3.75 .723 .700 .460
Number of Semantic Sets 17.00 16.50 16.75 .662 .472 .312 9.25 10.25 8.25 .675 .646 .382
Semantic Entropy 4.00 5.25 4.75 .792 .613 .525 13.25 13.25 13.25 .637 .623 .278
Perplexity 12.00 10.75 11.75 .745 .558 .432 16.50 15.75 16.25 .603 .603 .225
TokenSAR 13.75 13.50 13.75 .735 .540 .413 14.25 13.25 14.50 .616 .615 .243
SentenceSAR 9.00 10.25 9.00 .773 .571 .483 13.75 15.00 14.25 .631 .602 .263
MaxNLI 20.25 20.25 20.25 .466 .303 -.036 20.00 19.00 20.00 .514 .532 .035
AlignScore 17.00 18.25 17.50 .631 .415 .207 7.50 7.50 7.25 .682 .666 .372
Parametric Knowledge 18.00 18.00 17.75 .627 .425 .247 18.25 19.25 18.75 .561 .556 .104
XGBoost (all UQ features) 9.50 8.50 7.25 .766 .594 .494 3.00 3.25 3.75 .733 .705 .462
XGBoost (franq features) 13.75 14.75 14.00 .723 .526 .409 9.25 7.25 8.00 .672 .670 .368
franq no calibration 14.25 12.25 14.75 .728 .553 .403 11.00 10.00 10.75 .664 .641 .345
franq calibrated 2.75 4.00 2.75 .797 .628 .537 7.00 7.75 6.00 .696 .672 .411
franq condition-calibrated 2.75 3.00 2.25 .802 .631 .541 2.50 3.00 2.75 .730 .711 .477
Table 3: Results averaged across four QA datasets for Llama 3B and Falcon 3B. For each model, the first three columns show the mean rank across all 4 datasets, and the last three columns report the mean metric values. Lower mean ranks and higher mean values indicate better performance. The condition-calibrated franq is top-performing across all settings, except mean ROC-AUC on Falcon 3B, where it ranks second.

### v3

@@ 2.3 Calibrating franq
Since the UQ methods $\text{UQ}_{\text{faith}}$ and $\text{UQ}_{\text{unfaith}}$ of equation (4) may have different distributions, to avoid inconsistencies and miscalibration among various UQ measures, we calibrate their outputs using isotonic regression on the training data Vashurin et al. (2025).
Formally, given training dataset $\mathcal{D}=\{(u_{i},\text{fact}_{i})\}_{i=1}^{N}$ comprising pairs of UQ scores $u_{i}$ and corresponding binary factuality labels $\text{fact}_{i}$ for the $N$ claims, we calibrate the UQ scores by fitting a non-decreasing function $f\colon\mathbb{R}\rightarrow[0,1]$ through isotonic regression, minimizing the squared error:
@@
where $\mathcal{F}$ denotes the set of all non-decreasing functions mapping real numbers to probabilities in the interval $[0,1]$. Isotonic regression directly optimizes over $\mathcal{F}$ without assuming any parametric or functional form for $f$. This yields a piecewise-constant function $\widehat{f}$ defined on the observed UQ scores that satisfies the monotonicity constraint. During inference, we apply the calibration function $\widehat{f}$ to each UQ score to produce probabilistically meaningful output.
Condition-Calibrated franq. Since $\text{UQ}_{\text{faith}}$ and $\text{UQ}_{\text{unfaith}}$ represent factuality scores under faithful and unfaithful conditions, respectively, we propose condition-specific calibration. This involves partitioning the training dataset $\mathcal{D}$ into two subsets: faithful claims $\mathcal{D}_{\text{faith}}$ and unfaithful claims $\mathcal{D}_{\text{unfaith}}$. Then, we calibrate $\text{UQ}_{\text{faith}}$ using the subset $\mathcal{D}_{\text{faith}}$ and $\text{UQ}_{\text{unfaith}}$ using the subset $\mathcal{D}_{\text{unfaith}}$.
We consider franq with condition-specific calibration as our primary method. To evaluate the impact of calibration, we additionally assess two variants: one without any calibration, and another one in which both UQ methods are calibrated using the full training dataset $\mathcal{D}$. The calibration strategies are summarized as follows:
@@
    Calibrated. Both UQ methods are calibrated on the entire training dataset $\mathcal{D}$, disregarding claim faithfulness.
@@
    Condition-calibrated. Each UQ method is calibrated using a subset of the training data corresponding to the respective condition: $\text{UQ}_{\text{faith}}$ is calibrated using $\mathcal{D}_{\text{faith}}$, and $\text{UQ}_{\text{unfaith}}$ is calibrated using $\mathcal{D}_{\text{unfaith}}$.
@@ 3 Datasets for RAG Uncertainty Quantification
Existing datasets for studying RAG hallucinations have serious limitations, as they typically evaluate only context-relative correctness rather than factuality (see Section 5). We argue that factuality is more critical in RAG applications, with faithfulness serving as a complementary perspective. Consequently, an effective dataset should capture both factual errors and contextual misuse.
To address this need, we introduce a new dataset specifically designed for long-form generations, enabling fine-grained analysis of atomic claims.
3.1 Long-form QA Dataset.
Questions. Our long-form QA dataset consists of 76 questions: 44 most challenging questions from RAGTruth Niu et al. (2024) (identified by those with highest number of hallucinated claims), and 32 additional technical “how-to” questions generated using GPT-4 via simple prompts (e.g., requesting challenging, domain-diverse technical questions such as “How does solar power generate electricity?”). The generated questions were manually inspected to ensure clarity and relevance.
Retrieval Model. For each question, we retrieve the top-k=3 passages using the Facebook Contriever model Izacard et al. (2021) with embeddings computed over the 2018 English Wikipedia, ensuring high-quality and reliable evidence passages.
LLMs. We construct four model-specific dataset subsets by generating long-form answers to all 76 questions with their corresponding retrieved passages, using greedy decoding independently for each model: Llama 3B Instruct, Llama 8B Instruct Grattafiori et al. (2024), Falcon 3B Base Team (2024), and Gemma 4B Instruct Team et al. (2025). These subsets enable UQ methods to be evaluated on top of in-policy generations for each model.
@@
Claim Extraction. For each generated answer, we extract atomic claims and their corresponding token spans using the approach of Wang et al. (2024b); Vashurin et al. (2025). First, GPT-4o extracts decontextualized atomic claims from the entire text paragraph through a dedicated prompt. Then, for each claim, a second prompt instructs GPT-4o to list the relevant words from the original text, which we map to token spans. Applying this procedure, we obtain 1,782 claims for Llama 3B Instruct and 1,548 claims for Falcon 3B Base. From these claims, we select 500 claims for train set, reserving the remainder for test set. The prompts used for claims extraction and mapping are listed in Appendix B.
@@
Annotation. We annotate factuality and faithfulness using GPT-4o-search with dedicated prompts. Faithfulness labels categorize each claim as either faithful or unfaithful. Factuality annotations include three categories: True, False, and Unverifiable. We retain only verifiable claims (True or False), binarizing the labels accordingly.
During verification of the automatic annotations, we found that the False and Unverifiable categories are particularly difficult to assess automatically, so we manually reviewed all claims assigned to either category and corrected their labels when necessary.
Further details on prompts, dataset statistics and annotation scheme are provided in Appendix B.
3.2 Short-form QA Datasets
In contrast to long-form QA, where evaluating factuality requires extracting model-specific claims and annotating them, short-form QA provides gold-standard answers for each question. This allows us to directly compare each model’s generated answer with the ground-truth answer, yielding an automatic factuality judgment without additional manual annotation or claim-level verification.
@@ 3.2 Short-form QA Datasets
Questions. We adapt four short-form QA datasets for RAG evaluation: TriviaQA Joshi et al. (2017), SimpleQA Wei et al. (2024a), Natural Questions Kwiatkowski et al. (2019), and PopQA Mallen et al. (2023). For each dataset, we sample 200 questions for training and 1000 for testing, and we treat each model response as a single claim.
RAG Models. We use the same retrieval model as in the long-form setting, selecting the top-k=5 passages per question. For LLMs, we use the same two Llama models and the Falcon model, along with an additional model: Gemma 12B Instruct Team et al. (2025).
Annotation. We evaluate factuality of each generated answer by comparing it against the gold-standard answer using GPT-4o, following the procedure of Wei et al. (2024a), which has been shown to yield reliable factuality judgments.
@@ 4 Experiments
In this section, we evaluate franq and corresponding baselines on both the short-form and long-form benchmarks described in Section 3. For all experiments, we fix the retrieval process and the underlying white-box LLM, and we assess the factual accuracy of the model-generated claims.
Later, through ablation studies, we examine the contribution of the individual franq components, $P(\text{faithful})$, $\text{UQ}_{\text{faith}}$, and $\text{UQ}_{\text{unfaith}}$, as well as the effect of varying the amount of training data.
@@ 4.1 Experimental Setup
Method Llama 3B Instruct Falcon 3B Base Llama 8B Instruct Gemma 4B Instruct
PR-AUC $\uparrow$ PRR $\uparrow$ PR-AUC $\uparrow$ PRR $\uparrow$ PR-AUC $\uparrow$ PRR $\uparrow$ PR-AUC $\uparrow$ PRR $\uparrow$
General Baselines
Max Claim Prob. .058 -.029 .126 .258 .055 .118 .061 .0
P(True) .117 .207 .077 .170 .071 .112 .096 .148
Perplexity .056 -.081 .090 .165 .075 .090 .048 -.071
Max Token Entropy .109 .115 .130 .219 .102 .138 .051 -.003
CCP .085 .169 .162 .181 .061 .108 .087 .216
RAG-Specific Baselines
AlignScore .075 .108 .104 .233 .068 .119 .061 .058
Parametric Knowledge .064 .018 .067 .029 .059 .047 .112 .183
XGBoost
XGBoost (all UQ features) .124 .206 .088 .198 .044 -.015 .073 .085
XGBoost (franq features) .111 .149 .080 .086 .048 .017 .090 .158
franq
franq no calibration .100 .181 .135 .362 .063 .162 .080 .200
franq calibrated .103 .256 .074 .090 .043 -.047 .150 .401
franq condition-calibrated .140 .223 .173 .354 .081 .184 .090 .208
@@
UQ baselines. We group all UQ methods into four categories: (1) general baselines, (2) RAG-specific baselines, (3) XGBoost-based methods, and (4) three variants of our proposed franq method, each using a different calibration strategy.
General baselines. We compare franq with general baselines, which consist of standard UQ methods applied directly to the LLM’s output distribution without using any RAG-specific structure. For implementation, we use the LM-Polygraph library Fadeeva et al. (2023). A complete list of methods we used is provided in Table 1.
RAG-specific baselines. We also evaluate the two franq components in isolation, AlignScore and Parametric Knowledge, to assess how much their combination in franq improves over using each component individually (see Section 2.2).
XGBoost methods. We include XGBoost models trained on factuality labels using two feature sets: (1) the three components used in franq (AlignScore, $\text{UQ}_{\text{faith}}$, $\text{UQ}_{\text{unfaith}}$), and (2) all available unsupervised UQ method.
FRANQ. Finally, we evaluate three franq variants with different calibration strategies for $\text{UQ}_{\text{faith}}$ and $\text{UQ}_{\text{unfaith}}$ (see Section 2.3): no calibration, calibrated, and condition-calibrated.
Evaluation measures. Each UQ method produces factuality estimates, which we compare against binary gold-standard labels using PR-AUC, treating false claims as the positive class to emphasize their detection. We also assess rejection performance using the Prediction Rejection Ratio (PRR; Mallen et al., 2023) with a maximum rejection threshold of 0.5. PRR measures how effectively the model rejects uncertain predictions while retaining accurate ones, capturing its ability to prioritize reliable outputs.
@@ 4.2 Long-Form QA Results
For long-form QA, we evaluate each UQ method using PR-AUC and PRR across four models (Llama 3B Instruct, Falcon 3B Base, Llama 8B Instruct and Gemma 4B Instruct), see Table 2. The condition-calibrated franq achieves the best PR-AUC and second-best PRR for Llama 3B Instruct and Falcon 3B Base, while for Llama 8B Instruct it attains the best PRR and second-best PR-AUC. The calibrated franq achieves the highest PRR for Llama 3B Instruct and the highest PR-AUC and PRR for Gemma 4B Instruct. The non-calibrated franq also performs strongly, ranking first and second in PRR for Falcon 3B Base and Llama 8B Instruct, respectively. Overall, franq demonstrates strong and consistent performance across all models.
Method Llama 3B Instruct Falcon 3B Base Llama 8B Instruct Gemma 12B Instruct
PR-AUC $\uparrow$ PRR $\uparrow$ PR-AUC $\uparrow$ PRR $\uparrow$ PR-AUC $\uparrow$ PRR $\uparrow$ PR-AUC $\uparrow$ PRR $\uparrow$
General Baselines
Max Sequence Prob. .558 .454 .628 .256 .569 .407 .400 .162
Mean Token Entropy .594 .481 .613 .242 .640 .491 .423 .230
CCP .551 .443 .641 .304 .553 .417 .412 .198
Lexical Similarity .564 .479 .618 .277 .639 .532 .430 .240
Degree Matrix .629 .520 .702 .464 .627 .492 .464 .260
Sum of Eigenvalues .628 .518 .700 .460 .628 .489 .467 .260
Semantic Entropy .613 .525 .623 .278 .637 .519 .466 .261
SentenceSAR .571 .483 .602 .263 .556 .414 .416 .174
RAG-specific Baselines
AlignScore .415 .207 .666 .372 .432 .224 .376 .158
Parametric Knowledge .425 .247 .556 .104 .499 .330 .364 .105
XGBoost
XGBoost (all UQ features) .594 .494 .705 .462 .634 .503 .474 .301
XGBoost (franq features) .526 .409 .670 .368 .524 .385 .414 .196
franq
franq no calibration .553 .403 .641 .345 .523 .340 .447 .225
franq calibrated .628 .537 .672 .411 .644 .534 .481 .258
franq condition-calibrated .631 .541 .711 .477 .647 .540 .496 .283
[…]